15 women can knit 6 sweaters in 20 days working 6 hours a day. Find in how many days can 20 women knit 10 sweaters working 3 hours a day?
Answer by richwmiller(17219)

15*20*6=1800 hours for 6 sweaters or 300 hours per sweater.
300*10=3000 hours for 10 sweaters.
20*3=60 hours
3000/60=50 days for 10 sweaters by 20 women working 3 hours a day.